5. Advertising expenses are a significant component of the cost of products sold. The following distribution of frequencies in a sample of the advertising expenses of 70 manufacturing companies. The standard deviation is:

Advertising expenses (in thousands of $)

Number of companies

25 to 35

36 to 46

47 to 57

58 to 68

69 to 79

5

10

21

16

8

a. $11,123

b. $12,452

c. $13,345

d. $14,467

e. $15,588
Answer #1
Expenditure Mid-Point (x) No of Companies (f) xf (x - x̄) (x - x̄)^2 f(x - x̄)^2
25 - 35 30 5 150 -24.2 585.64 2,928.2
36 - 46 41 10 410 -13.2 174.24 1742.4
47 - 57 52 21 1092 -2.2 4.84 101.64
58 - 68 63 16 1008 8.8 77.44 1239.04
69 - 79 74 8 592 19.8 392.04 3136.32
Total 60 3252 9147.6

Mean = Σxf / Σf = 3252 / 60 = 54.2

Standard Deviation S = √Σf(x - x̄)^2 / n - 1

= √9147.6 / 60 - 1

= √155.044068

= 12.452